# https://ssugarwh.github.io/consistentid.github.io/

### v1

@@ Abstract
Diffusion-based technologies have made significant strides, particularly in personalized and customized facial generation. However, existing methods face challenges in achieving high-fidelity and detailed identity (ID) consistency, primarily due to insufficient fine-grained control over facial areas and the lack of a comprehensive strategy for ID preservation by fully considering intricate facial details and the overall face. To address these limitations, we introduce ConsistentID, an innovative method crafted for diverse identity-preserving portrait generation under fine-grained multimodal facial prompts, utilizing only a single reference image. ConsistentID comprises two key components: a multimodal facial prompt generator that combines facial features, corresponding facial descriptions and the overall facial context to enhance precision in facial details, and an ID-preservation network optimized through the facial attention localization strategy, aimed at preserving ID consistency in facial regions. Together, these components significantly enhance the accuracy of ID preservation by introducing fine-grained multimodal ID information from facial regions. To facilitate training of ConsistentID, we present a fine-grained portrait dataset, FGID, with over 500,000 facial images, offering greater diversity and comprehensiveness than existing public facial datasets. Experimental results substantiate that our ConsistentID achieves exceptional precision and diversity in personalized facial generation, surpassing existing methods in the MyStyle dataset. Furthermore, while ConsistentID introduces more multimodal ID information, it maintains a fast inference speed during generation. Our codes and pre-trained checkpoints will be available at https://github.com/JackAILab/ConsistentID.
Keywords:
Portrait generation fine-grained conditions identity preservation
Xiaodan Liang is the corresponding author.
@@
1 Introduction
Recently, image-generation technology  [16, 25, 41, 14, 46] has undergone significant evolution, driven by the emergence and advancement of diffusion-based [45, 11] text-to-image large models like GLIDE [30], DALL-E 2 [36], Imagen[42], Stable Diffusion (SD) [37], eDiff-I [1] and RAPHAEL [55]. This progress has given rise to a multitude of application approaches across diverse scenarios. Positioned as the central focus of these application approaches, personalized and customized portrait generation has attracted widespread attention in both academic and industrial domains, owing to its extensive applicability in downstream tasks such as E-commerce advertising, personalized gift customization and virtual try-ones.
To improve the precision and diversity of personalized portrait generation with reference images, numerous customized methodologies have emerged, meeting users’ demands for high-quality customized images. These personalized approaches are categorized based on whether fine-tuning occurs during inference, resulting in two distinct types: test-time fine-tuning and direct inference. Test-time fine-tuning: This category includes methods such as Textual Inversion [8], HyperDreambooth[41], and CustomDiffusion[20]. Users can achieve personalized generation by providing a set of target ID images for post-training. Despite achieving commendable high-fidelity results, the quality of the generated output depends on the quality of manually collected data. Additionally, the manual collection of customized data for fine-tuning introduces a labor-intensive and time-consuming aspect, limiting its practicality. Direct inference: Another category of models, including IP-Adapter [56], Fastcomposer[54], Photomaker[22], and InstantID[53], adopts a single-stage inference approach. These models enhance global ID consistency by either utilizing the image as a conditional input or manipulating image-trigger words. However, most methods frequently overlook fine-grained information, such as landmarks and facial features. Although InstantID improves ID consistency to some extent by introducing landmarks, the visual prompt landmark restricts the diversity and variability of key facial regions, leading to stiff generated facial features. In summary, two pivotal challenges requiring meticulous consideration persist in personalized portrait generation: 1) neglect of fine-grained facial information and 2) identity inconsistency between facial areas and the whole face, as illustrated in Figure 5.
To address these challenges, we introduce a novel method, ConsistentID, crafted to maintain identity consistency and capture diverse facial details through multimodal fine-grained ID information, employing only a single facial image while ensuring high fidelity. Figure 2 provides the overview of our ConsistentID. ConsistentID comprises two key modules: 1) a multimodal facial prompt generator and 2) an ID-preservation network. The former component includes a fine-grained multimodal feature extractor and a facial ID feature extractor, enabling the generation of more detailed facial ID features using multi-conditions, incorporating facial images, facial regions, and their corresponding textual descriptions extracted from the multimodal large language model LLaVA1.5[23]. Utilizing the facial ID features obtained from the initial module, we feed them into the latter module, promoting ID consistency across each facial region via the facial attention localization strategy. Additionally, we recognize the limitations of existing portrait datasets[61, 3, 27, 31, 52], particularly in capturing diverse and fine-grained identity-preserving facial details, crucial for the effectiveness of ConsistentID. To address this, we introduce the inaugural Fine-Grained ID Preservation (FGID) dataset, along with a fine-grained identity consistency metric, providing a unique and comprehensive evaluation approach to enhance our training and performance evaluation in facial details.
@@ 1 Introduction
    We introduce ConsistentID to improve fine-grained customized facial generation by incorporating detailed descriptions of facial regions and local facial features. Experimental results showcase the superiority of ConsistentID in terms of ID consistency and high fidelity, even with just one reference image. Simultaneously, despite the introduction of more detailed multimodal fine-grained ID information in ConsistentID, the inference speed remains relatively efficient, as shown in Table  1
@@ 2 Related Work
Text-to-image Diffusion Models. Diffusion models have made notable advancements, garnering significant attention from both industry and academia, primarily due to their exceptional semantic precision and high fidelity. The success of these models can be attributed to the utilization of high-quality image-text datasets, continual refinement of foundation modality encoders, and the iterative enhancement of controlled modules. In the domain of text-to-image generation, the encoding of text prompts involves utilizing a pretrained language encoder, such as CLIP[35], to transform it into a latent representation, subsequently inserted into the diffusion model through the cross-attention mechanism. Pioneering models in this realm encompass GLIDE [30], SD [37], DiT [32], among others, with further developments and innovations continuing to emerge [13, 55]. A notable advancement in this lineage is SDXL [33], which stands out as the most powerful text-to-image generation model. It incorporates a larger Unet [38] model and employs two text encoders for enhanced semantic control and refinement. As a follow-up, we use SD [37] model as our base model to achieve personalized portrait generation.
@@ 3.1 Multimodal Facial Prompt Generator
1) Text Embedding. Motivated by recent works[26, 25, 29, 50] in personalized facial generation, our goal is to introduce more detailed and accurate facial descriptions. To achieve this, we input the entire facial image into the Multimodal Large Language Model (MLLM) LLaVA1.5 [23] using the command prompt ‘Describe this person’s facial features, including face, ears, eyes, nose, and mouth’ to obtain a description at the facial feature level. Subsequently, we replace the words ‘face, ears, eyes, nose, and mouth’ in these feature-level descriptions with the delimiter ‘<facial>’ and concatenate them with the captions of the entire facial image. Finally, the concatenated descriptions are fed into the pre-trained text encoder to learn fine-grained multimodal facial features. With both visual and textual descriptions containing more precise ID information, our ConsistentID effectively mitigates ID inconsistency issues in facial details.
@@
2) Facial Embedding. In contrast to existing methods [28, 17], [28, 17, 8, 9, 24, 39, 54, 22, 53]. that rely on brief textual descriptions or coarse-grained visual prompts, our goal is to integrate more fine-grained multimodal control information at the facial region level. This aims to achieve improved accuracy in identity (ID) consistency and diverse facial generation. To enrich the ID-preservation information, we delve into more fine-grained facial features, including eye gaze, earlobe characteristics, nose shape, and others. Following the previous method [15, 48, 58, 21, 59, 18],we employ the pre-trained face model BiSeNet [57] to extract segmentation masks of facial areas, encompassing eyes, nose, ears, mouth, and other regions, from the entire face. Subsequently, the facial region images obtained from these masks are fed into the pre-trained image encoder to learn fine-grained facial embeddings. The inclusion of facial regions’ features results in fine-grained facial embeddings containing more abundant ID-preservation information compared to features learned from the entire face. More detailed processes are outlined in the supplementary materials.
@@
3) Facial Encoder. Previous studies [35, 54, 8, 40, 57] have demonstrated that relying solely on visual or textual prompts cannot comprehensively maintain ID consistency both in appearance and semantic details. While IP-Adapter makes the initial attempt to simultaneously inject multimodal information through two distinct decoupled cross-attention mechanisms, it overlooks ID information from crucial facial regions, rendering it susceptible to ID inconsistency in facial details.
@@
To cultivate the potential of image and text prompts, inspired by the token fusion approach of multimodal large language models, we design a facial encoder to seamlessly integrate visual prompts with text prompts along the dimension of the text sequence, as depicted in Figure 3. Specifically, given a facial embedding and a caption embedding, the facial encoder initially employs a self-attention mechanism to align the entire facial features with facial areas’ features, resulting in aligned features denoted as ${\widehat{f}^{i}\in\mathbb{R}^{N\times D}}$, where $N=5$ represents the number of facial feature areas, including eyes, mouth, ears, nose, and other facial regions, and $D$ represents the dimension of text embeddings. In cases where face images lack a complete set of N facial features, the missing features are padded using an all-zero matrix. Subsequently, we replace the text features at the position of the delimiter ‘<facial>’ with $\widehat{f}^{i}$, and then employ two multi-layer perceptron (MLP) to learn the text conditional embeddings.
@@
Facial ID Feature Extractor. Except for the input condition of fine-grained facial features, we also inject the character’s overall ID information into our ConsistentID as a visual prompt. This process relies on the pre-trained CLIP image encoder and the pre-trained face model from the specialized version of the IP-Adapter [56] model, IPA-FaceID-Plus [56]. Specifically, the complete facial images are concurrently fed into both encoders for visual feature extraction. After these two encoders, a lightweight projection module, with parameters initialized by IPA-FaceID-Plus, is used to generate the face embedding of the whole image.
@@ 3.2 ID-Preservation network
The effectiveness of image prompts in pre-trained text-to-image diffusion models [54, 22, 60, 4, 34, 47] significantly enhances textual prompts, especially for content that is challenging to describe textually. However, visual prompts alone often provide only coarse-grained control due to the semantic fuzziness of visual tokens. To solve this, we follow the IP-Adapter and integrate fine-grained multimodal ID prompts and overall ID prompts into the UNet model through the cross-attention module to achieve precise ID preservation.
@@
Motivated by Fastcomposer [54], we introduce an ID-consistent network to maintain consistency in local ID features by directing the attention of facial features to align with the corresponding facial regions. This optimization strategy is derived from the observation that traditional cross-attention maps tend to simultaneously focus on the entire image, posing a challenge in maintaining ID features during the generation of facial regions. To address this issue, we introduce facial segmentation masks during training to obtain attention scores learned from the enhanced text cross-attention module for facial regions.
@@
Let $P\in[0,1]^{h\times w\times n}$ represent the cross-attention map that connects latent pixels to multimodal conditional embeddings at each layer, where $P[i,j,k]$ signifies the attention map from the $k$-th conditional token to the ($i$, $j$) latent pixel. Ideally, the attention maps of facial region tokens should focus exclusively on facial feature areas, preventing the blending of identities between facial features and averting propagation to the entire facial image. To achieve this goal, we propose localizing the cross-attention map using the segmentation mask of reference facial regional features.
@@
Let $M=\{m_{1},m_{2},m_{3},...,m_{N}\}$ represent the segmentation masks of the reference portrait, $I=\{i_{1},i_{2},i_{3},...,i_{N}\}$ as the index list indicating which facial feature corresponds to visual and textual tokens in the multimodal prompt, and $P_{i}=P[:,:,i]\in[0,1]^{h\times w}$ denote the cross-attention map of the $i$-th facial region’s token, where $I$ is generated using the special token ‘<facial>’. Given the cross-attention map $P_{i_{j}}$, it should closely correspond to the facial region identified by the $j$-th multimodal token, segmented by $m_{j}$. To achieve this, we introduce $m_{j}$ and apply it to $P_{i_{j}}$ to obtain its corresponding activation region, aligning with the segmentation mask $m_{j}$ of the $j$-th facial feature token. For achieving this correspondence, a balanced L1 loss is employed to minimize the distance between the cross-attention map and the segmentation mask:
@@
$\mathcal{L}_{\text{loc }}=\frac{1}{N}\sum_{j=1}^{N}\left(\operatorname{mean}\left(P_{i_{j}}\left[1-m_{j}\right]\right)-\operatorname{mean}\left(P_{i_{j}}\left[m_{j}\right]\right)\right),$ (1)
@@
where N denotes the number of the segmentaion masks. This loss formulation aims to ensure that each facial feature token’s attention map aligns closely with its corresponding segmentation mask, promoting precise and localized attention during the generation process.

### v2

# ConsistentID[Uncaptioned image]: Portrait Generation with Multimodal Fine-Grained Identity Preserving

@@ Abstract
Diffusion-based technologies have made significant strides, particularly in personalized and customized facial generation. However, existing methods struggle to achieve high-fidelity and detailed identity (ID) consistency. This is mainly due to two challenges: insufficient fine-grained control over specific facial areas and the absence of a comprehensive strategy for ID preservation that accounts for both intricate facial details and the overall facial structure. To address these limitations, we introduce ConsistentID, an innovative method crafted for diverse identity-preserving portrait generation under fine-grained multimodal facial prompts, utilizing only a single reference image. ConsistentID comprises two core components: a multimodal facial prompt generator and an ID-preservation network. The facial prompt generator combines localized facial features, facial feature descriptions, and overall facial descriptions to enhance the precision of facial detail reconstruction. The ID-preservation network, optimized with a facial attention localization strategy, ensures consistent identity preservation across facial regions. Together, these components leverage fine-grained multimodal identity information to improve identity preservation accuracy significantly. To drive ConsistentID’s training, we propose a fine-grained portrait dataset, FGID, with over 500,000 facial images, offering greater diversity and comprehensiveness than existing public facial datasets. Experimental results substantiate that our ConsistentID achieves exceptional precision and diversity in personalized facial generation, surpassing existing methods in the MyStyle dataset. In addition, although ConsistentID introduces more multimodal ID information, it still maintains rapid inference speed during the generation process. Our codes and pre-trained checkpoints are available at https://github.com/JackAILab/ConsistentID.
Index Terms:
Portrait generation, fine-grained control, identity preservation.
1 Introduction
I-mage-generation technology [1, 2, 3, 4, 5] has undergone significant evolution, driven by the emergence and advancement of diffusion-based [6, 7] text-to-image large models like GLIDE [8], DALL-E 2 [9], Imagen [10], Stable Diffusion (SD) [11], eDiff-I [12] and RAPHAEL [13]. This progress has given rise to a multitude of application approaches across diverse scenarios. Among these, personalized portrait generation has garnered significant attention in both academia and industry due to its wide-ranging applications, including e-commerce advertising, personalized gift customization, and virtual try-ons.
To improve the precision and diversity of personalized portrait generation with reference images, numerous customized methodologies have emerged, meeting users’ demands for high-quality customized images. These personalized approaches are categorized based on whether fine-tuning occurs during inference, resulting in two distinct types: test-time fine-tuning and direct inference. Test-time fine-tuning: This category includes methods such as Textual Inversion [18], HyperDreambooth[3], and CustomDiffusion[19]. Users can achieve personalized generation by providing a set of target ID images for post-training. Despite achieving commendable high-fidelity results, the quality of the generated output depends on the quality of manually collected data. Additionally, the manual collection of customized data for fine-tuning introduces a labor-intensive and time-consuming aspect, limiting its practicality. Direct inference: Another category of models, including IP-Adapter [15], Fastcomposer[20], Photomaker [21], and InstantID [22], adopts a single-stage inference approach. These models enhance global ID consistency by either utilizing the image as a conditional input or manipulating image-trigger words. However, most methods frequently overlook fine-grained information, such as landmarks and facial features. Although InstantID [22] enhances ID consistency to some extent by incorporating landmarks, the use of visual prompt landmarks limits the diversity and flexibility of key facial regions, resulting in rigidly generated facial features. Building on these observations, it becomes evident that two pivotal challenges persist in personalized portrait generation, requiring meticulous consideration: 1) neglect of fine-grained facial information and 2) identity inconsistency between facial areas and the whole face, as illustrated in Figure 9.
@@ 1 Introduction
To tackle identity consistency and detail preservation challenges in personalized image generation, we propose ConsistentID, a novel method that excels in maintaining identity fidelity while capturing diverse facial details. With just a single facial image as input, our approach leverages multimodal fine-grained ID features to deliver high-quality results. As illustrated in Figure 1, ConsistentID enables personalized transformations such as revitalizing old photos, altering gender, or changing outfits. By seamlessly integrating facial inputs with text prompts, our model generates diverse yet consistent identity representations, offering unmatched adaptability in personalized generation tasks.
Figure 2 provides an overview of our ConsistentID. ConsistentID comprises two key modules: 1) a multimodal facial prompt generator and 2) an ID-preservation network. The former component includes a fine-grained multimodal feature extractor and a facial ID feature extractor, enabling the generation of more detailed facial ID features using multi-conditions, incorporating facial images, facial regions, and corresponding textual descriptions extracted from the multimodal large language model LLaVA1.5 [23]. The facial ID features generated by the first module are then passed to the latter module, which enhances ID consistency across facial regions through a facial attention localization strategy. Additionally, we recognize the limitations of existing portrait datasets [24, 25, 26, 27, 28], particularly in capturing diverse and fine-grained facial details that preserve identity, crucial to the effectiveness of ConsistentID. To address this, we introduce the inaugural Fine-Grained ID Preservation (FGID) dataset, along with a fine-grained identity consistency metric, providing a unique and comprehensive evaluation approach to enhance our training and provide a comprehensive evaluation of its performance in capturing facial details.
@@
    We introduce ConsistentID to improve fine-grained customized facial generation by incorporating detailed descriptions of facial regions and local facial features. The experimental results demonstrate the superiority of ConsistentID in terms of ID consistency and high fidelity, even with only one reference image. Additionally, although ConsistentID introduced more detailed multimodal fine-grained ID information during training, it achieves this with a single fixed prompt  [^1] This person has one nose, two eyes, two ears, and a mouth. during inference. It does not depend on facial descriptions generated by LLaVA1.5, ensuring a streamlined and efficient approach, as shown in Table III.
@@ 2 Related Work
Text-to-image Diffusion Models. Diffusion models have made notable advancements, garnering significant attention from both industry and academia, primarily due to their exceptional semantic precision and high fidelity. The success of these models can be attributed to the utilization of high-quality image-text datasets, continual refinement of foundation modality encoders, and the iterative enhancement of controlled modules. In the domain of text-to-image generation, the encoding of text prompts involves utilizing a pretrained language encoder, such as CLIP [29], to transform it into a latent representation, subsequently inserted into the diffusion model through the cross-attention mechanism. Pioneering models in this domain include GLIDE [8], SD [11], DiT [30], among others, and further developments and innovations are continuing to emerge [31, 13]. A notable advancement in this lineage is SDXL [32], which stands out as the most powerful text-to-image generation model. It incorporates a larger Unet [33] model and employs two text encoders for enhanced semantic control and refinement. As a follow-up, we use the SD [11] model as our base model to achieve personalized portrait generation.
@@ 3 Method
In this section, we will introduce our multimodal facial prompt generator in Subsection 3.1 and propose a meticulously designed ID-preservation network in Subsection 3.2. Next, we detail the training process of ConsistentID and its inference procedure in Subsection 3.3. Finally, we provide a comprehensive description of the FGID dataset in Subsection 3.4.
@@ 3.1 Multimodal Facial Prompt Generator
1) Text Embedding. Motivated by recent works [36, 2, 37, 38] in personalized facial generation, our goal is to introduce more detailed and accurate facial descriptions. To achieve this, the entire facial image is processed by the Multimodal Large Language Model (MLLM) LLaVA1.5 [23] using the prompt: ‘Describe this person’s facial features, including the face, ears, eyes, nose, and mouth’. The mentioned process will generate feature-level descriptions of the facial areas. Next, the terms ‘face’, ‘ears’, ‘eyes’, ‘nose’, and ‘mouth’ in these descriptions are replaced with the delimiter ‘$<$facial$>$’, and the modified text is concatenated with captions describing the entire facial image. Finally, the concatenated descriptions are input into a pre-trained text encoder to learn the fine-grained text embedding of the facial features. With the abundant descriptions from facial regions, the text embeddings are enriched with more precise identity information, effectively resolving identity inconsistency.
@@
2) Facial Embedding. In contrast to existing methods [39, 40], [39, 40, 18, 41, 42, 43, 20, 21, 22] that rely on brief textual descriptions or coarse-grained visual prompts, our goal is to integrate more fine-grained multimodal control information at the facial region level, aiming to achieve greater accuracy in capturing detailed facial features. To enrich the ID-preservation information, we delve into more fine-grained facial features, including eye gaze, earlobe characteristics, nose shape, and others. Following the previous method [44, 45, 46, 47, 48, 49], we employ the pre-trained face model BiSeNet [50] to extract segmentation masks of facial areas, encompassing eyes, nose, ears, mouth, and other regions, from the entire face. Subsequently, the facial regions obtained from these masks are fed into the pre-trained image encoder to learn fine-grained facial embeddings. The inclusion of facial regions’ features results in fine-grained facial embeddings containing more abundant ID-preservation information compared to features learned from the entire face.
@@
3) Facial Encoder. Previous studies [29, 20, 18, 34, 50] have demonstrated that relying solely on visual or textual prompts cannot comprehensively maintain ID consistency both in appearance and semantic details. While IP-Adapter [15] makes the initial attempt to simultaneously inject multimodal information through two distinct decoupled cross-attention mechanisms, it overlooks ID information from crucial facial regions, rendering it susceptible to ID inconsistency in facial details.
@@
To cultivate the potential of image and text prompts, inspired by the token fusion approach of multimodal large language models, we design a facial encoder to seamlessly integrate visual prompts with text prompts along the dimension of the text sequence, as depicted in Figure 3. Specifically, given a facial embedding and a caption embedding, the facial encoder initially employs a self-attention mechanism to align the entire facial features with facial areas’ features, resulting in aligned features denoted as ${\widehat{f}^{i}\in\mathbb{R}^{N\times D}}$, where $N=5$ represents the number of facial feature areas, including eyes, mouth, ears, nose, and other facial regions, and $D$ represents the dimension of text embeddings. In cases where face images lack a complete set of $N$ facial features, the missing features are padded using an all-zero matrix. Subsequently, we replace the text features at the position of the delimiter ‘$<$facial$>$’ with $\widehat{f}^{i}$ using the visual token replacement operation, as illustrated in Figure 3 (right). Finally, the text features, now enriched with visual identity information, are fed into two multi-layer perceptions (MLP) to learn the text conditional embeddings.
@@
Facial ID Feature Extractor. Except for the input condition of fine-grained facial features, we also inject the character’s overall ID information into our ConsistentID as a visual prompt. This process relies on the pre-trained CLIP image encoder and the pre-trained face model from the specialized version of the IP-Adapter [15] model, IPA-FaceID-Plus [15]. Specifically, the complete facial images are simultaneously fed into both encoders for visual feature extraction. Following these two encoders, a lightweight projection module, with parameters initialized by IPA-FaceID-Plus, is used to generate the face embedding of the whole image.
@@ 3.2 ID-Preservation Network
The effectiveness of visual prompts in pre-trained text-to-image diffusion models [20, 21, 14, 51, 52, 53] significantly enhances textual prompts, especially for content that is challenging to describe textually. However, visual prompts alone often provide only coarse-grained control due to the semantic fuzziness of visual tokens. To solve this, we integrate fine-grained multimodal ID prompts and overall ID prompts into the UNet model through the cross-attention module to achieve precise ID preservation.
Specifically, we first introduced an ID consistency network, which maintains the consistency of local ID features by guiding the attention of facial features to align with the corresponding facial regions [20].
@@
This optimization strategy is derived from the observation that traditional cross-attention maps tend to simultaneously focus on the entire image rather than specific facial regions, making it challenging to maintain ID features during each facial region generation. To address this issue, we introduce facial segmentation masks during training to obtain attention scores learned from the enhanced text cross-attention module for facial regions.
@@
Let $P\in[0,1]^{h\times w\times n}$ represent the cross-attention map that connects latent pixels to multimodal conditional embeddings at each layer, where $P[i,j,k]$ signifies the attention map from the $k$-th conditional token to the ($i$, $j$) latent pixel. Ideally, the attention maps of facial region tokens should focus exclusively on facial feature areas, preventing the blending of identities between facial features and averting propagation to the entire facial image. To achieve this goal, we propose localizing the focus of the cross-attention map by using segmentation masks that correspond to the reference facial regional features.
@@
Let $M=\{m_{1},m_{2},m_{3},...,m_{N}\}$ represent the segmentation masks of the reference facial regions, $I=\{i_{1},i_{2},i_{3},...,i_{N}\}$ as the index list indicating which facial feature corresponds to visual and textual tokens in the multimodal prompt, and $P_{i}=P[:,:,i]\in[0,1]^{h\times w}$ denote the cross-attention map of the $i$-th facial region’s token, where $I$ is generated using the special token ‘$<$facial$>$’.
Given the cross-attention map $P_{i_{j}}$, it should closely correspond to the facial region identified by the $j$-th multimodal token, segmented by $m_{j}$. To achieve this, we introduce $m_{j}$ and apply it to $P_{i_{j}}$ to obtain its corresponding activation region, aligning with the segmentation mask $m_{j}$ of the $j$-th facial feature token. For achieving this correspondence, a balanced $\mathcal{L}_{1}$ loss is employed to minimize the distance between the cross-attention map and the segmentation mask:
@@
$\mathcal{L}_{\text{facial}}=\frac{1}{N}\sum_{j=1}^{N}\left(\operatorname{mean}\left(P_{i_{j}}\left[1-m_{j}\right]\right)-\operatorname{mean}\left(P_{i_{j}}\left[m_{j}\right]\right)\right),$ (1)
@@
where N denotes the number of the segmentation masks. This loss formulation aims to ensure that each facial feature token’s attention map aligns closely with its corresponding segmentation mask, promoting precise and localized attention during the generation process.